# Deep-ELA: Deep Exploratory Landscape Analysis with Self-Supervised Pretrained Transformers for Single- and Multi-Objective Continuous Optimization Problems

### v1

@@ 3.1 Model Structure
Figure 4: Illustration of the chosen topology of the backbone model without the training heads. The model receives $(\mathcal{X},\mathcal{Y})$ as input and outputs $\mathcal{F}_{\text{D-ELA}}$ – and, optionally, the embedding of the tokens $\mathcal{T}_{\text{Final}}$ after the final LayerNorm. $\mathcal{T}_{\text{Final}}$ is only relevant for the contrastive loss during training and is ignored after training. The initial $k$NN embedding is used to capture the local information of all points from the input sample, and followed by a stride operator to optionally reduce the number of tokens without losing information. Next, the model consists of six Multi-Head Attention blocks, followed by a Feed-Forward block of two successive Linear layers each. The LayerNorm layers are after the shortcuts as proposed by Nguyen and Salazar, (2019). We chose GLU activations in the Feed Forward layers with a 4×4$\times$4 × larger number of hidden neurons. The last Linear + GLU layer projects the high-dimensional embeddings into lower dimensions. Afterward, the mean over all tokens is computed and normalized into $[-1,1]$ by a Tanh activation.
@@
Following this linear projection, the sequence of tokens is sequentially processed by six combined modules of Multi-Head Attention (MHA) and Feed-Forward (FF), concluding with a final Layer Normalization (LN; Ba et al.,, 2016). Through experimentation, we determined that six iterations strike an optimal balance between model intricacy and performance. The output post-LN layer manifests as a normalized $(n\times d_{\text{model}})$-dimensional tensor, with $n$ representing the number of tokens. To derive the features $\mathcal{F}_{\text{D-ELA}}$, the final embedding is condensed via a linear layer with GLU activation, referred to as feature-extractor, into a $\mathbb{R}^{n\times{c_{\text{D-ELA}}}}$ tensor. Ultimately, a mean pooling step across the token sequence yields ${c_{\text{D-ELA}}}$ features, which are constrained within $[-1,1]$ by a concluding Tanh activation. An illustration of the model’s topology and an overview of its associated hyperparameters are given in Figure 4 and Table 1, respectively.
This architecture termed the backbone model, is versatile for diverse downstream tasks. However, as the backbone essentially functions as a feature generator, it lacks trainability. Consequently, to facilitate training, we incorporate (1.) two feature-heads (distinguishing one as a so-called student and the other one as a teacher), and (2.) an additional teacher feature-extractor, as visualized in Figure 5. The student components are updated via gradient descent, while the teacher counterparts are updated via exponential moving average (EMA). Both are integral for the self-supervised loss computation. The training loss strategy revolves around providing distinct, augmented versions of the same objective instance to both the teacher and student. Here, the teacher generates target projections from which the student gleans insights. However, we refrained from using two backbone models (one for the student and one for the teacher) as this would drastically increase the total amount of model parameters. So, the backbone model is the same for both the student and teacher. More details about the loss function and the training routine are provided in Chapter 3.1.2.
@@ 3.1.1 $k$NN Embedding
$\mathbf{x}_{1}$$\mathbf{x}_{4}$$\mathbf{x}_{2}$$\mathbf{x}_{5}$$\mathbf{x}_{3}$ind $k$NN of $\mathbf{x}_{i}\in X$ and create tokens $\mathbf{t}_{i}\in T$$\overbrace{\phantom{xxxxx}}^{\textbf{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}Global Context}}$$\mathbf{t}_{1}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{1},\mathbf{y}_{1}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{4}-\mathbf{x}_{1}),(\mathbf{y}_{4}-\mathbf{y}_{1}),(\mathbf{x}_{2}-\mathbf{x}_{1}),(\mathbf{y}_{2}-\mathbf{y}_{1})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{2}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{2},\mathbf{y}_{2}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{1}-\mathbf{x}_{2}),(\mathbf{y}_{1}-\mathbf{y}_{2}),(\mathbf{x}_{4}-\mathbf{x}_{2}),(\mathbf{y}_{4}-\mathbf{y}_{2})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{3}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{3},\mathbf{y}_{3}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{4}-\mathbf{x}_{3}),(\mathbf{y}_{4}-\mathbf{y}_{3}),(\mathbf{x}_{5}-\mathbf{x}_{3}),(\mathbf{y}_{5}-\mathbf{y}_{3})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{4}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{4},\mathbf{y}_{4}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{1}-\mathbf{x}_{4}),(\mathbf{y}_{1}-\mathbf{y}_{4}),(\mathbf{x}_{2}-\mathbf{x}_{4}),(\mathbf{y}_{2}-\mathbf{y}_{4})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{5}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{5},\mathbf{y}_{5}},\underbrace{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}(\mathbf{x}_{2}-\mathbf{x}_{5}),(\mathbf{y}_{2}-\mathbf{y}_{5}),(\mathbf{x}_{1}-\mathbf{x}_{5}),(\mathbf{y}_{1}-\mathbf{y}_{5})\ldots}}_{\textbf{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\pgfsys@color@rgb@stroke{1}{.5}{0}\pgfsys@color@rgb@fill{1}{.5}{0}Local Context}}\big{)}^{\text{T}}$
@@ 3.1.2 Contrastive Loss
In this equation, $H$ and $D_{\text{KL}}$ denote the entropy and the Kullback-Leibler divergence, respectively. The cross-entropy approach provides pronounced gradients, even for vanishing activations, compared to the e.g. mean squared error (MSE). The loss function also depends on the critical hyperparameter $\tau$, which is the temperature of the Softmax activation $\sigma$. Usually, the value of $\tau$ is set in the range $(0,0.3]$. The smaller $\tau$ is, the higher the loss for hard negative samples, i.e., distinct instances that yield similar projections and, thus, are hard for the model to differentiate. On the contrary, for larger values of $\tau$, the loss is more evenly spread over hard and easy negative samples. Through initial testing, we found that $\tau=0.05$ works well in our scenario. To increase the likelihood of the appearance of hard negative samples within a single batch, the batch size is supposed to be as large as possible, which is why we aimed for very large training batches.
@@ 4.1 Black-Box Optimization Problems
Most validation and testing optimization problems come from BBOB, provided by the COmparing Continuous Optimisers (COCO; Hansen et al.,, 2019) platform. COCO offers a diverse range of continuous optimization problem suites, spanning from single- and multi-objective to noisy and mixed-integer problems. This study primarily engaged with the 24 noiseless, single-objective functions from the (classical) BBOB suite. For each of these functions, 1 00010001\,0001 000 unique instances were produced, of which the first 500 served for testing and the latter for validation purposes (see Chapter 5). Function dimensionalities were chosen from $d\in\{2,3,5,10\}$. Although the suite includes functions with $d\in\{20,40\}$, our model’s current dimensional constraints prevent their inclusion.
For the multi-objective case study, we analyzed the same multi-objective problems that Rook et al., (2022) examined. These bi-objective ($m=2$) functions have a decision space dimensionality of $d=2$. Specifically, we considered the Bi-BBOB functions $f_{46},f_{47},$ and $f_{50}$, along with the ZDT, DTLZ, and MMF function groups (excluding ZDT5 and MMF13) from the smoof package. Further details are given in Chapter 5.4.
@@ 4.2 Random Optimization Problems
Refer to caption (a) Dispersion
Refer to caption (b) Meta
Refer to caption (c) $Y$-Distribution
@@
Tian et al., (2020) introduced a method capable of generating an arbitrary number of optimization problems in milliseconds. This approach hinges on a random tree-like structure constituted of various operators such as mean, sum, exp, log, among many others. This generator operates on three input parameters: the number of dimensions of the decision space ($d$), and the lower and upper bounds of the number of operators. Once initiated, the generator fabricates an optimization problem for a $d$-dimensional set of decision variables $\mathcal{X}$ by randomly assorting a number of operators bounded by the provided lower and upper bounds. The result is a randomly constructed optimization problem that yields the objective values $Y$. We manually calibrated the generator’s hyperparameters to craft optimization problems that are similar to BBOB. Additionally, we incorporated extra operators to amplify the linear and squared dependencies among decision and objective variables, noting that the original generator somewhat diverged from the BBOB suite in these aspects. Furthermore, we tweaked the skewness and kurtosis of the objective variables for closer alignment with BBOB (refer to Figure 7).
@@ 5.1 Pretraining
We chose to train both a medium- and a large-sized model. The medium model can manage a total dimensionality of six, while the large model can accommodate up to twelve total dimensions. The large model has twice the width of the medium model and produces twice the number of features (48 vs. 24). Due to the squared growth of transformer complexity with the number of tokens, we introduced a stride factor. A stride of two implies that every second token is omitted after the $k$NN-embedding, as depicted in Figure 4. Consequently, the large model’s complexity is reduced to a fourth, compared to a model with a stride of one, maintaining a larger batch size during the training of the large model. We then trained each model-variant with sample sizes of $25d$ and $50d$, resulting in four final models. For the $50d$ models, we reduced the batch size by half but compensated with gradient accumulation over two iterations. However, InfoNCE does not benefit much from gradient accumulation since its efficiency relies on a substantial number of samples. Still, we opted for gradient accumulation to maintain a consistent number of update steps and training instances, as seen with the $25d$ models.
For training, we employed PyTorch-Lightning [^3] https://lightning.ai/ in tandem with PyTorch [^4] https://pytorch.org/. Post-training, we removed the two heads necessary during the training phase, resulting in a versatile backbone model suitable for various downstream tasks. Subsequently, we incorporated classical machine learners atop this backbone. No further fine-tuning of the backbone model was undertaken or deemed necessary. However, it is possible to fine-tune the backbone model for different tasks, although this demands a deep understanding of deep learning techniques. To emulate a user proficient in general machine learning but not deeply versed in deep learning specifics, we relied exclusively on Scikit-Learn and refrained from fine-tuning the backbone model. Further, we did not apply feature selection to demonstrate that the instance features generated by the Deep-ELA approach contain features of high relevance and little redundancy.

### v2

@@ 3.2 Model Structure
Figure 4: Illustration of the chosen topology of the backbone model without the training heads. The model receives $(\mathcal{X},\mathcal{Y})$ as input and outputs $\mathcal{F}_{\text{D-ELA}}$ – and, optionally, the embedding of the tokens $\mathcal{T}_{\text{Final}}$ after the final LayerNorm. $\mathcal{T}_{\text{Final}}$ is only relevant for the contrastive loss during training and is ignored after training. The initial $k$NN embedding (Seiler et al.,, 2022) is used to capture the local information of all points from the input sample, and followed by a stride operator to optionally reduce the number of tokens without losing information. Next, the model consists of six Multi-Head Attention blocks, followed by a Feed-Forward block of two successive Linear layers each. The LayerNorm layers are after the shortcuts as proposed by Nguyen and Salazar, (2019). We chose GLU activations in the Feed Forward layers with a 4×4$\times$4 × larger number of hidden neurons. The last Linear + GLU layer projects the high-dimensional embeddings into lower dimensions. Afterward, the mean over all tokens is computed and normalized into $[-1,1]$ by a Tanh activation.
@@
Following this linear projection, the sequence of tokens is sequentially processed by six combined modules of Multi-Head Attention (MHA) and Feed-Forward (FF), concluding with a final Layer Normalization (LN; Ba et al.,, 2016). Through experimentation, we determined that six iterations strike an optimal balance between model intricacy and performance. This choice was also supported by Vaswani et al., (2017) as they also opted for six iterations. The output post-LN layer manifests as a normalized $(n\times d_{\text{model}})$-dimensional tensor, with $n$ representing the number of tokens. To derive the features $\mathcal{F}_{\text{D-ELA}}$, the final embedding is condensed via a linear layer with GLU activation, referred to as feature-extractor, into a $\mathbb{R}^{n\times{c_{\text{D-ELA}}}}$ tensor. Ultimately, a mean pooling step across the token sequence yields ${c_{\text{D-ELA}}}$ features, which are constrained within $[-1,1]$ by a concluding Tanh activation. An illustration of the model’s topology and an overview of its associated hyperparameters are given in Figure 4 and Table 1, respectively.
This architecture termed the backbone model, is versatile for diverse downstream tasks. However, as the backbone essentially functions as a feature generator, it lacks trainability. Consequently, to facilitate training, we incorporate (1.) two feature-heads (distinguishing one as a so-called student and the other one as a teacher), and (2.) an additional teacher feature-extractor, as visualized in Figure 5. The student components are updated via gradient descent, while the teacher counterparts are updated via exponential moving average (EMA). Both are integral for the self-supervised loss computation. The training loss strategy revolves around providing distinct, augmented versions of the same objective instance to both the teacher and student. Here, the teacher generates target projections from which the student gleans insights. However, we refrained from using two backbone models (one for the student and one for the teacher) as this would drastically increase the total amount of model parameters. So, the backbone model is the same for both the student and teacher. More details about the loss function and the training routine are provided in Section 3.2.2.
@@ 3.2.1 $k$NN Embedding
$\mathbf{x}_{1}$$\mathbf{x}_{4}$$\mathbf{x}_{2}$$\mathbf{x}_{5}$$\mathbf{x}_{3}$ind $k$NN of $\mathbf{x}_{i}\in X$ and create tokens $\mathbf{t}_{i}\in T$$\overbrace{\phantom{xxxxx}}^{\textbf{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}Global Context}}$$\mathbf{t}_{1}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{1},\mathbf{y}_{1}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(\mathbf{x}_{4}-\mathbf{x}_{1}),(\mathbf{y}_{4}-\mathbf{y}_{1}),(\mathbf{x}_{2}-\mathbf{x}_{1}),(\mathbf{y}_{2}-\mathbf{y}_{1})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{2}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{2},\mathbf{y}_{2}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(\mathbf{x}_{1}-\mathbf{x}_{2}),(\mathbf{y}_{1}-\mathbf{y}_{2}),(\mathbf{x}_{4}-\mathbf{x}_{2}),(\mathbf{y}_{4}-\mathbf{y}_{2})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{3}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{3},\mathbf{y}_{3}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(\mathbf{x}_{4}-\mathbf{x}_{3}),(\mathbf{y}_{4}-\mathbf{y}_{3}),(\mathbf{x}_{5}-\mathbf{x}_{3}),(\mathbf{y}_{5}-\mathbf{y}_{3})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{4}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{4},\mathbf{y}_{4}},{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(\mathbf{x}_{1}-\mathbf{x}_{4}),(\mathbf{y}_{1}-\mathbf{y}_{4}),(\mathbf{x}_{2}-\mathbf{x}_{4}),(\mathbf{y}_{2}-\mathbf{y}_{4})\ldots}\big{)}^{\text{T}}$$\mathbf{t}_{5}=\big{(}{\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\mathbf{x}_{5},\mathbf{y}_{5}},\underbrace{{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}(\mathbf{x}_{2}-\mathbf{x}_{5}),(\mathbf{y}_{2}-\mathbf{y}_{5}),(\mathbf{x}_{1}-\mathbf{x}_{5}),(\mathbf{y}_{1}-\mathbf{y}_{5})\ldots}}_{\textbf{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}Local Context}}\big{)}^{\text{T}}$
@@ 3.2.2 Contrastive Loss
In this equation, $H$ and $D_{\text{KL}}$ denote the entropy and the Kullback-Leibler divergence, respectively. The cross-entropy approach provides pronounced gradients, even for vanishing activations, compared to the e.g. mean squared error (MSE). The loss function also depends on the critical hyperparameter $\tau$, which is the temperature of the Softmax activation $\sigma$. Usually, the value of $\tau$ is set in the range $(0,0.3]$. The smaller $\tau$ is, the higher the loss for hard negative samples, i.e., distinct instances that yield similar projections and, thus, are hard for the model to differentiate. On the contrary, for larger values of $\tau$, the loss is more evenly spread over hard and easy negative samples. Through initial testing, we found that $\tau=0.05$ works well in our scenario. As an explanation, we argue that small $\tau$-values work better in our scenario as we rely solely on the stochasticity of the synthetic data generator. We cannot control the characteristics of the generated instances. So it is very likely that most instances are easy to distinguish and hard negatives occur rather infrequently. To account for this, we opted for two solutions: (1) a small $\tau$-value will penalize hard negatives effectively also in large batches and (2) utilizing large batches to increase the likelihood of the appearance of hard negative samples within a single batch.
@@ 4.1 Black-Box Optimization Problems
Most validation and testing optimization problems come from BBOB, provided by the COmparing Continuous Optimisers (COCO; Hansen et al.,, 2019) platform. COCO offers a diverse range of continuous optimization problem suites, spanning from single- and multi-objective to noisy and mixed-integer problems. This study primarily engaged with the 24 noiseless, single-objective functions from the (classical) BBOB suite. For each of these functions, 1 00010001\,0001 000 unique instances were produced, of which the first 500 served for testing and the latter for validation purposes (see Section 5). Function dimensionalities were chosen from $d\in\{2,3,5,10\}$. Although the suite includes functions with $d\in\{20,40\}$, our model’s current dimensional constraints prevent their inclusion.
For the multi-objective case study, we analyzed the same multi-objective problems that Rook et al., (2022) examined. These bi-objective ($m=2$) functions have a decision space dimensionality of $d=2$. Specifically, we considered the Bi-BBOB functions $f_{46},f_{47},$ and $f_{50}$, along with the ZDT, DTLZ, and MMF function groups (excluding ZDT5 and MMF13) from the smoof package. Further details are given in Section 5.4.
@@ 4.2 Random Optimization Problems
(a) Dispersion
(b) Meta
(c) $Y$-Distribution
@@
Tian et al., (2020) introduced a method capable of generating an arbitrary number of optimization problems in milliseconds. This approach hinges on a random tree-like structure constituted of various operators such as mean, sum, exp, log, among many others. This generator operates on three input parameters: the number of dimensions of the decision space ($d$), and the lower and upper bounds of the number of operators. Once initiated, the generator fabricates an optimization problem for a $d$-dimensional set of decision variables $\mathcal{X}$ by randomly assorting a number of operators bounded by the provided lower and upper bounds. The result is a randomly constructed optimization problem that yields the objective values $Y$. We manually calibrated the upper and lower bounds of the generator to craft optimization problems that are similar to BBOB. We found that when choosing these bounds too small or too large, the generator produces objectives that are too simple or too complicated in comparison to BBOB. Yet, as the models (like any other machine learner) can only generalize to instances that are similar to the training data, we need to ensure that BBOB is not out-of-distribution to the models. We found that upper and lower bounds of $(4,32)$ fulfill this criterion well. Additionally, we incorporated extra operators to amplify the linear and squared dependencies among decision and objective variables, noting that the original generator somewhat diverged from the BBOB suite in these aspects. Furthermore, we tweaked the skewness and kurtosis of the objective variables for closer alignment with BBOB (refer to Figure 7). Last, the original Python implementation yielded NaN-values quite often. We identified and modified those operators that cause the appearance of NaN-values. Afterward, we observed that the generator produces diverse optimization problems that are more similar to BBOB in terms of their complexity of ELA features (see Figure 7). Yet, it shall be noted that the random functions are not identical to BBOB. Instead, the random functions cover similar characteristics as BBOB.
@@ 5.1 Pretraining
We chose to train both a medium- and a large-sized model. The medium model can manage a total dimensionality of six, while the large model can accommodate up to twelve total dimensions. The large model has twice the width of the medium model and produces twice the number of features (48 vs. 24). Due to the squared growth of transformer complexity with the number of tokens, we introduced a stride factor. A stride of two implies that every second token is omitted after the $k$NN-embedding, as depicted in Figure 4. Consequently, the large model’s complexity is reduced to a fourth, compared to a model with a stride of one, maintaining a larger batch size during the training of the large model. We then trained each model-variant with sample sizes of $25d$ and $50d$, resulting in four final models, where $d$ stands for the dimensions of $\mathcal{X}$. Hence, the sample-size scales linearly with the number of dimensions of the decision space. For the $50d$ models, we reduced the batch size by half but compensated with gradient accumulation over two iterations. However, InfoNCE does not benefit much from gradient accumulation since its efficiency relies on a substantial number of samples. Still, we opted for gradient accumulation to maintain a consistent number of update steps and training instances, as seen with the $25d$ models.
For training, we employed PyTorch-Lightning [^4] https://lightning.ai/ in tandem with PyTorch [^5] https://pytorch.org/. Post-training, we removed the two heads necessary during the training phase, resulting in a versatile backbone model suitable for various downstream tasks. Subsequently, we incorporated classical machine learners atop this backbone. No further fine-tuning of the backbone model was undertaken or deemed necessary. Instead, the backbone model can be directly applied as the considered studies have similar boxed constraints for $\mathcal{X}$ as the training instances. Yet, if e.g. the box constraints are noticeably dissimilar to the training set ($[-5,5]$) fine-tuning may be advisable. But this requires further testing in future work. To emulate a user proficient in general machine learning but not deeply versed in deep learning specifics, we relied exclusively on Scikit-Learn and refrained from fine-tuning the backbone model. Further, we did not apply feature selection to demonstrate that the instance features generated by the Deep-ELA approach contain features of high relevance and little redundancy. Last, it should be noted that the backbone model did not receive any instances from the BBOB suite or similar frameworks and was solely trained on randomly generated instances. Further, the model receives each optimization instance exactly once. Hence, we argue that any overfitting to existing optimization problems is hardly possible. More importantly, as the models were not trained to predict certain labels as it is usually done in supervised learning, they cannot just remember labels for a specific task.